Can you please help me with something? My question is how do i solve proportions?
I've been trying to figure that out and i was hoping you could help me. Thank you for your time to help me. You are nice

Answer by Alan3354(69443)

Can you please help me with something? My question is how do i solve proportions?
------------
If you're given:
a:b::c:d (a is to b as c is to d)
The product of the means (inside terms) = product of the extremes (outside terms)
------------
a*d = b*c
That's all.
------------
It's the same as
a/b = c/d